A standing wave on a string of length L = 2 m has a wavelength λ = 0.5 m. The strings oscillates in its:

1
Expert's answer
2021-12-20T10:29:32-0500

The wavelength of a standing wave in a string is given as follows


"\lambda = \dfrac{2L}{n}"


where "n" is the number of harmonic. Expressing "n", get:



"n = \dfrac{2L}{\lambda} = \dfrac{2\cdot 2m}{0.5m} = 8"

Answer. The strings oscillates in its 8th harmonic.